The first overtone of an open pipe has frequency n. The first overtone of closed pipe of same lenght will have frequency A)n/2B)2nC)4nD)n/4

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
for open organ pipe fundamental frequency is f1 = v/2l.
for closed organ pipe fundamental frequency is f2 = v/4l.
f1/f2 = 2
Hence
Frequency of closed organ pipes = n/2
Regards
Arun (askIITians forum expert)